# Self-Tuning Spectral Clustering for Speaker Diarization

### v1

@@ III Experimental Setup
We empirically compare the unsupervised speaker diarization performances of the proposed SC-pNA (with $p=20\%$ retention), with the EER-$\Delta$ approach, and the SOTA method of ASC. As a baseline, the semi-supervised speaker diarization performance of CSC is also measured. We also empirically validate the robustness of SC-pNA by observing its performance over a range of retention percentages $p$ from $10$ to $50$, with increments of $5$.
@@ III-A Dataset
For our experiments, we utilized the speech corpora from the third DIHARD speech diarization challenge (DIHARD-III) [6]. It comprises data from eleven diverse domains characterized by a variety in the number of speakers, conversation types and speech qualities. For this study, we focused on seven domains: broadcast interview, court, cts, maptask, meeting, socio lab, and webvideo. These domains are representative of everyday conversational speech scenarios between adults. We excluded four domains because they either contained single-speaker data (audiobooks) or regions with personally identifiable information (PII) (clinical, restaurant, and socio field data). For each domain there exists two splits of development (dev) and evaluation (eval). Both splits are used to evaluate the SD methods, as the methods operate in an unsupervised manner. This results in a total of fourteen data splits over the seven data domains.
@@ III-B SD system
We conducted our experiments using the SpeechBrain toolkit [24]. Our implementation is based on a modified version of the AMI recipe provided in the toolkit  [^1] https://github.com/speechbrain/speechbrain/tree/develop/recipes/AMI, which employs a pre-trained ECAPA-TDNN model as the speaker embedding extractor [25]. The ECAPA-TDNN model was trained on the VoxCeleb dataset [26, 27]. The extracted speaker embeddings are 192-dimensional, derived from the penultimate layer of the ECAPA-TDNN architecture. To ensure robust performance and eliminate the variability introduced by SAD, we utilized ground truth annotations. Each segment is set to a duration of 3.0 seconds, with an overlap of 1.5 seconds between consecutive segments.
@@ III-C Evaluation metric
We use standard DER mertric for evaluating the performance of the diarization system [28]. DER is comprised of three key errors: missed speech, false alarm of speech, and speaker error. DER is quantified as the ratio of the combined duration of these three errors to the total duration.
@@ IV Results & Discussion
IV-A Comparison with SOTA
The unsupervised SD performances of the proposed SC-pNA (with $p=20\%$ retention) and the EER-$\Delta$ approach are compared with the SOTA method of ASC. The resulting DERs achieved over the seven domains of DIHARD-III are shown in Table I. We observe that SC-pNA achieves the lowest DER among the unsupervised methods over nine of the fourteen data splits. ASC achieved the lowest DER among the remaining four data splits, whereas EER-$\Delta$ in general obtained higher DERs across the data splits. With the exception of the court domain, SC-pNA has in general obtained better DER compared to ASC, and is competitive on the two splits where ASC has lower DER. Thus we can recommend SC-pNA for the task of unsupervised SD.
TABLE I: Comparison of DER (lower is better) achieved by the unsupervised SD methods of ASC, EER-$\Delta$, and SC-pNA, over seven domains of DIHARD-III. The CSC provides a semi-supervised SD baseline.
@@ IV-A Comparison with SOTA
Comparing the unsupervised SD performance of SC-pNA with the semi-supervised SD performance of CSC, we observe that with the exception of the court domain, SC-pNA in fact obtains DERs that are close to the DERs of CSC. For six of the data splits: cts (dev), maptask (eval), socio lab (dev and eval), and webvideo (dev and eval), the unsupervised SC-pNA even outperforms the semi-supervised CSC. Thus in general, the SD performance of SC-pNA is observed exceed the performance of the SOTA ASC, and is also competitive with respect to the semi-supervised method of CSC.
@@ IV-B Variations in DERs for different $p\%$ in SC-pNA
The previous study involved SC-pNA with retention percentage $p$ as $20\%$. Here we study the variation in the resulting DER for values of $p$ in the range of $10\%$ to $50\%$, with an increment of $5\%$. For four data splits of maptask (eval), meeting (eval), socio lab(eval), and webvideo (eval), in Figure 2 the DERs obtained by SC-pNA are shown, and are compared with those obtained by CSC and ASC. From this figure, we observe that for maptask (eval) and webvideo (eval), the DERs obtained across all retention percentages are better than not only those obtained by ASC, by also those obtained by the semi-supervised CSC. For meeting (eval) the DER is generally higher than ASC or CSC, and we observe that at the retention percentages of $25\%$ and $20\%$ the DERs are the closest to ASC. For socio lab (eval) the DERs are also lower than ASC, and quite comparable to the baseline CSC. Thus we conclude that the proposed SC-pNA in general shows robust SD performance across different retention percentages, and the selected $20\%$ retention can be recommended as it generally achieves low DERs across various data splits.
Figure 2: For four data splits, the DERs (lower is better) obtained by SC-pNA over different retention percentages are compared with the DERs obtained by CSC and ASC.
@@ V Conclusion
We proposed an SD approach called SC-pNA, which creates a sparse affinity matrix following an approach motivated by the theory on EER for Gaussian distributions. SC-pNA identifies two clusters in each row of an initial affinity matrix, and retains the top $p\%$ similarity scores from the cluster with larger similarities. This approach enables SC-pNA to prune a variable number of similarity scores. The resulting sparse affinity matrix is symmetrized, from which the Laplacian is formed. The speakers are then identified by clustering the spectral embeddings, based on the number of speakers that is automatically selected from the eigengap. The overall method of SC-pNA has significantly lower computation cost compared to the SOTA ASC approach, while our empirical results show that SC-pNA with a retention percentage of $20\%$ outperforms the ASC over the DIHARD-III dataset, making SC-pNA a method to be recommended for SD. Empirical results also notably showed that SC-pNA was competitive against the semi-supervised CSC, and showed robust performances across different retention percentages.
@@ Acknowledgment
The primary author expresses sincere gratitude to the Linguistic Data Consortium (LDC) for the LDC Data Scholarship, which enabled access to the DIHARD-III dataset.

### v2

@@ III Experimental Setup
We empirically compare the unsupervised SD performances of the proposed SC-pNA (with $p=20\%$ retention), with the EER-$\Delta$, and the SOTA method of ASC. The performance of the semi-supervised CSC is also measured, which serves as an ideal baseline for unsupervised SD. For all methods, we set $k_{\max}$ to 10. We also empirically validate the robustness of SC-pNA by observing its performance over a range of retention percentages $p$ from $10$ to $50$, with increments of $5$.
@@ III-A Dataset
For our experiments, we utilized the speech corpora from the third DIHARD speech diarization challenge (DIHARD-III) [6]. It comprises data from eleven diverse domains characterized by a variety in the number of speakers, conversation types and speech qualities. For each domain there exists two splits of development (dev) and evaluation (eval). This results in a total of twenty-two data splits over the eleven data domains. Both splits are used to evaluate the unsupervised SD methods.
@@ III-B SD system
We conducted our experiments using the SpeechBrain toolkit [24]. Our implementation is based on a modified version of the AMI recipe provided in the toolkit  [^1] https://github.com/speechbrain/speechbrain/tree/develop/recipes/AMI, which extracts speaker embeddings using an ECAPA-TDNN model [25], pre-trained on the VoxCeleb dataset [26, 27]. The extracted speaker embeddings are 192-dimensional, derived from the penultimate layer of the ECAPA-TDNN architecture. To ensure robust performance and eliminate the variability introduced by SAD, we utilized ground truth annotations. Each segment is set to 3.0 seconds, with 1.5 seconds of overlap between consecutive segments.
@@ III-C Evaluation metric
We use the standard DER metric for evaluating the performance of the diarization system [28]. DER is comprised of three key errors: missed speech, false alarm of speech, and speaker confusion error. DER is quantified as the ratio of the combined duration of these three errors to the total duration.
@@ IV Results & Discussion
IV-A Comparison with the SOTA
The proposed SC-pNA (with $p=20\%$ retention) and the EER-$\Delta$ approach are compared with the SOTA method of ASC. The resulting DERs achieved over DIHARD-III are shown in Table I.
TABLE I: Speaker Diarization performance in terms of DER (lower is better) for the semi-supervised CSC baseline along with the unsupervised SD methods of ASC, EER-$\Delta$, and SC-pNA, on DIHARD-III dataset.
@@ IV-A Comparison with the SOTA
restaurant dev 29.08 30.46 52.78 37.31
eval 29.93 33.35 59.78 38.81
audiobooks dev 0.40 23.64 0.20 0.08
eval 0.50 27.53 0.18 0.09
clinical dev 7.33 11.83 28.07 6.31
eval 4.34 10.62 31.23 3.31
socio field dev 7.01 7.31 23.38 9.21
eval 5.18 10.30 16.57 3.79
Overall dev 10.76 13.26 21.10 10.70
eval 9.97 13.11 20.08 10.27
We observe that SC-pNA achieves the lowest DER among the unsupervised methods over fourteen of the twenty-two data splits. ASC achieved the lowest DER among the remaining seven data splits, whereas EER-$\Delta$ in general obtained higher DERs across the data splits. SC-pNA achieves lower DER than ASC in six domains, is competitive in three, and has higher DER in only two domains. Across DIHARD-III, SC-pNA clearly outperforms both ASC and EER-$\Delta$, with performance comparable with the semi-supervised CSC. SC-pNA even outperforms CSC on the data splits: cts (dev), maptask (eval), socio lab (dev and eval), webvideo (dev and eval), audiobooks (dev and eval), clinical (dev and eval), and socio field (eval). Thus SC-pNA is recommended for unsupervised SD.
@@ IV-B Variations in DERs for different $p\%$ in SC-pNA
The preceding experiments evaluated the performance of SC-pNA with $p=20\%$. Here we study the variations in DER for values of $p$ in the range of $10\%$ to $50\%$, with an increment of $5\%$. For four data splits of maptask (eval), meeting (eval), socio lab(eval), webvideo (eval), in Figure 2 the DERs obtained by SC-pNA are compared with those obtained by CSC and ASC. Remarkably, for maptask (eval) and webvideo (eval), the DERs obtained across all retention percentages are better than both ASC and the semi-supervised CSC. For meeting (eval) the DER is generally higher than ASC or CSC, and at retention percentages of $25\%$ and $20\%$ the DERs are the closest to ASC. For socio lab (eval) the DERs are also lower than ASC, and are actually comparable to CSC. Thus we conclude that the proposed SC-pNA in general shows robust SD performance across different retention percentages, and the selected $20\%$ retention can be recommended as it generally achieves low DERs across various data splits.
Figure 2: The figure shows the variation in the SD performance in terms of DER across four data splits, for different retention percentages obtained by SC-pNA, CSC and ASC.
@@ V Conclusion
We proposed an SD approach called SC-pNA, which creates a sparse affinity matrix following an approach motivated by the theory on EER for Gaussian distributions. SC-pNA identifies two clusters in each row of an initial affinity matrix, and retains the top $p\%$ similarity scores from the cluster containing larger similarity scores. This approach enables SC-pNA to prune a variable number of similarity scores. The resulting sparse affinity matrix is symmetrized, from which the Laplacian is formed. The speakers are then identified by clustering the spectral embeddings, based on the number of speakers that is automatically selected from the eigengap. The overall method of SC-pNA has significantly lower computation cost compared to the SOTA ASC approach, while our empirical results show that SC-pNA with a retention percentage of $20\%$ outperforms the ASC over the DIHARD-III dataset, making SC-pNA a method to be recommended for SD. Empirical results also notably showed that SC-pNA was competitive against the semi-supervised CSC, and showed robust performances across different retention percentages.
@@ Acknowledgment
The primary author would like to thank his doctoral advisory committee members at RKMVERI for their feedback on the initial study on speaker diarization. He also expresses sincere gratitude to the Linguistic Data Consortium (LDC) for awarding the LDC Data Scholarship, which enabled access to the DIHARD-III dataset.